OpenStudy (anonymous):

Solve n^2 - 3n = 0 by factoring. Select all possible answers. 0 No solution 3 Others that aren't listed here. Infinite solutions

OpenStudy (anonymous):

take n as comon then n(n-3)=0 gives n=0.3

OpenStudy (anonymous):

its 0,3 noot 0.3..

OpenStudy (anonymous):

Thank you so much!

OpenStudy (nurali):

n^2 - 3n = 0 n(n-3)=0 n=0 or n-3=0 n=0 or n=3

OpenStudy (anonymous):

yw..